OpenStudy (anonymous):

find an equation of the line through (-2,-6);parallel to 3x+2y=5

OpenStudy (aroub):

parallel=same slope y=mx+b, m=-3/2 , (-2,-6) -6=-3/2(-2)+b -6=3+b -6-3=b -9=b so the equation is y=-3/2x-9

OpenStudy (anonymous):

let required line is 3x+2y=k since line is passing through point (-2,-6) ..hence -6-12=k k=-18 hence required line will be 3x+2y=-18 3x+2y+18=0

OpenStudy (aroub):

or yes you can say this is an correct answer but if if you want in the simplest form and in the form y=mx+b then its y=-3/2x-9

OpenStudy (anonymous):

i still dont understand how to do it though, sorry math doesn't come easy

OpenStudy (aroub):

oh so you want steps okay sure!! :)

OpenStudy (aroub):

well i actually gave steps i will just explain every step i did ok?

OpenStudy (anonymous):

okay

OpenStudy (aroub):

ok, i guess you know that parallel lines have equal slopes and you are given an equation or a line which is 3x+2y=5 now put that in the form y=mx+b do get the slope.. (and remember m is the slope) 2y=-3x+5 y= -3/2x+5/2 so u have -3/2 as a slope.. now you have slope whats left? THE Y-INTERCEPT! you get it by the points you are given which is (-2,-6) now you all you have to is substitute these points (-2 is x and -6 is y ) in the form y=mx+b y=mx+b (y=-6 , x=-2 , m=-3/2) u have to find B -6=-3/2(-2)+b -6=3+b -6-3=b -9=b so we found the y-intercept which is B (-9) and we found the slope which is m (-3/2) now we can form an equation.. y=-3/2x-9 im sorryyyy... im just not that good in explaining.. i hope you get some of what i just wrote! :) tell me if you have an question.. will be my pleasure answering it!! =)

OpenStudy (anonymous):

thanks your awesome

OpenStudy (aroub):

yaay! jk , im gllaaddddd i could heelpp!!! =)